OpenStudy (anonymous):

log(10 x 3cuberoot of 10) how to simplify that without using a calculator? I need the fastest numerical steps.

OpenStudy (anonymous):

@Auxuris :3

OpenStudy (gorv):

is it ...\[\log(10*\sqrt[3]{10})\]

OpenStudy (gorv):

????

OpenStudy (anonymous):

yeahp ;p

OpenStudy (anonymous):

sorry but didn't know how to write that down xD

OpenStudy (gorv):

its okay...both term are inmultiplication so use property of log

OpenStudy (gorv):

\[\log(a*b)=\log a + \log b\]

OpenStudy (anonymous):

okay you can split them up into ____ + _____

OpenStudy (gorv):

a=10 \[b=\sqrt[3]{10}\]

OpenStudy (anonymous):

and since 3√10 is basically 10^1/3 right?

OpenStudy (gorv):

log 10=1 \[\sqrt[3]{10}=10^{1/3}\]

OpenStudy (gorv):

\[\log x^y=y *logx \]

OpenStudy (anonymous):

wait so you can split the log outside to.. log10 because a "log" is log ten right?

OpenStudy (gorv):

use this property

OpenStudy (anonymous):

o.o

OpenStudy (gorv):

bring the power in gront and log 10 =1 again

OpenStudy (anonymous):

yes ^^ and when you have the same right hand side, you can add the two logs tgt log10 + (1/3)log10 = ?

OpenStudy (anonymous):

OOOH log10 = 1 right?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

log10 = 1(10)^1/3

OpenStudy (anonymous):

uh.. well, you can't do it like that because you still have a power you need it to be like, (a)log10 which will give you (a)(1)

OpenStudy (anonymous):

well I can't use a blasted calculator with this, what is 10^1/3? xD

OpenStudy (anonymous):

really? I can't do that??

OpenStudy (anonymous):

as i mentioned, shift the 1/3 in front

OpenStudy (anonymous):

log10^a = alog10

OpenStudy (anonymous):

oh so just do (1/3)(10)?

OpenStudy (anonymous):



OpenStudy (anonymous):

argh I'm hopeless xD Q_Q

OpenStudy (anonymous):

nope okay d'you remember this rule?